OpenStudy (anonymous):

1 question, will fan!

OpenStudy (anonymous):

What is that one question...

OpenStudy (anonymous):

@satellite73 could you help again, i'm trying to get things for my study guide right.

OpenStudy (anonymous):

\[f(5)=5500\ln(9\times 5+4)=5500\ln(49)\] and a calculator

OpenStudy (anonymous):

in plain english, where you see a \(t\) replace it by a \(5\)

OpenStudy (anonymous):

i got 275000

OpenStudy (anonymous):

let me check, seems large

OpenStudy (anonymous):

5500 times 49

OpenStudy (anonymous):

oh no dear, the log of 49, not 49

OpenStudy (anonymous):

Thats what ive been doing wrong then

OpenStudy (anonymous):

you forgot the \(\ln(49)\) part

OpenStudy (anonymous):

i get \(21405\) rounded http://www.wolframalpha.com/input/?i=5500ln%2849%29

OpenStudy (anonymous):

that make more sense?

OpenStudy (anonymous):

Ok, so what exactly is In? thats whats throwing me off.

OpenStudy (anonymous):

\(\ln(x)\) is a function you plug in a number, get out a number you usually have to use a calculator to evaluate it, so if i want \(\ln(49)\) like in this problem, you need a calculator

OpenStudy (anonymous):

Ok, thanks!, i only have a few more, think you could help?

OpenStudy (anonymous):

it is the inverse function of the exponential function \(f(x)=e^x\) in other words if \[\ln(x)=y\iff x=e^y\] you should be able to switch back and forth easily just like \(\log_2(x)=y\iff x=2^y\) sure shoot

OpenStudy (anonymous):

i mean "sure, go ahead and ask, i will answer or help"

OpenStudy (anonymous):

to increase a number by \(11\%\) multiply it by \(1.11\) as \(11\%=.11\)

OpenStudy (anonymous):

if you start with \(330\) deer, after 1 year there will be \[330\times 1.11\] deer after two year there will be \[330\times 1.11\times 1.11=330(1.11)^2\] deer, and after \(t\) years there will be \[330\times (1.11)^t\] deer

OpenStudy (anonymous):

ok, following so far.

OpenStudy (anonymous):

\[\large \color{red}{330}\times( \color{blue}{1.11})^{\color{green}t}\] \(\color{red}{330}\) is the original population of deer, what you have when you start counting

OpenStudy (anonymous):

\(\color{blue}{1.11}\) is \(1+.11\) how you increase a number by \(11\%\) and \(\color{green}t\) is the number of years after you start counting

OpenStudy (anonymous):

and so after 5 years you expect to have \[330\times (1.11)^5\] deer you need a calculator to compute this number

OpenStudy (anonymous):

i got 606.620936

OpenStudy (anonymous):

now not to confuse you, but if it is \(11\%\) compounded continuously and not "per year" which is what it says, then you would use \[\large 330\times e^{.11t}\]

OpenStudy (anonymous):

i get \(556\) http://www.wolframalpha.com/input/?i=330%281.11%29^5

OpenStudy (anonymous):

maybe you made a typo when you put it in your calculator i like wolfram cause you just type it in as you see it

OpenStudy (anonymous):

I am going to have to start using that! haha

OpenStudy (anonymous):

yeah, good idea

OpenStudy (anonymous):

Ok so after 5 years it will be 556?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

ok two more questions and the study guide is complete! :)

OpenStudy (anonymous):

the original price of the boat was $3500 so we know it will be \[3500\times b^x\] we just don't know \(b\) yet

OpenStudy (anonymous):

ok so start off with 3500 times b^x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now lets see if we can find another good point on the graph it looks like after 2 years it is at $2000

OpenStudy (anonymous):

so it has decreased by a factor of \(\frac{2000}{3500}=\frac{4}{7}\) so that will be our \(b\) but don't forget that was two year not one year

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so unless you want to estimate the value at \(t=1\) i would use \[3500\times \left(\frac{4}{7}\right)^{\frac{t}{2}}\]

OpenStudy (anonymous):

ok so right now b=4/7 and x=t/2?

OpenStudy (anonymous):

the reason for the 2 in the denominator of the exponent is because that was the decrease for two years, not one

OpenStudy (anonymous):

right, what you said

OpenStudy (anonymous):

to finish, compute at \(t=9.5\) and get \[3500\times \left(\frac{4}{7}\right)^{\frac{9.5}{2}}\]

OpenStudy (anonymous):

careful putting this in the calculator, there are lots of ways to make mistakes here

OpenStudy (anonymous):

im going to use that one website you were using, hold on.

OpenStudy (anonymous):

i got 8.59363

OpenStudy (anonymous):

ok if you use a calculator i would first compute \(9.5\div 2=4.75\) and only then compute \[3500(\frac{4}{7})^{4.75}\]

OpenStudy (anonymous):

careful with the syntax send me the link of what you wrote in wolfram

OpenStudy (anonymous):

i am saying that because i got something else

OpenStudy (anonymous):

something went wrong though.

OpenStudy (anonymous):

yeah what went wrong is you didn't put parentheses around the exponent you can see underneath what wolfram is computing, and it is not what you want

OpenStudy (anonymous):

that is why i mentioned computing \(9.5\div2=4.75\) first, then you wont have that mistake

OpenStudy (anonymous):

245.266

OpenStudy (anonymous):

much better

OpenStudy (anonymous):

Thats the final answer?

OpenStudy (anonymous):

since it is money, i would round

OpenStudy (anonymous):

\(\$245.27\) maybe

OpenStudy (anonymous):

Ok thats what i was typing haha

OpenStudy (anonymous):

or just \(\$245\) depending on how picky you want to be

OpenStudy (anonymous):

I'll just do 245.27

OpenStudy (anonymous):

k good that it?

OpenStudy (anonymous):

One last one, i think i know this one but im going to get your thoughts on it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

nice softball question there

OpenStudy (anonymous):

K is what controls to line rising and decreasing, when k is negative the line will decrease but when k is positive it will rise, since k is positive it is going to rise

OpenStudy (anonymous):

control the line vertically*

OpenStudy (anonymous):

okay slow here it is not a "line" but rather a "curve"

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

it shift the graph of \[y=ab^{x-h}+k\] up \(k\) units if \(k>0\)and down \(k\) units if \(k<0\) unless you are told \(k\) is positive, you cannot assume it is, because it is a variable

OpenStudy (anonymous):

i guess i should really say "it shifts the graph of \(y=ab^{x-h}\) up or down

OpenStudy (anonymous):

Alright, so k shifts the graph of y=ab^x-h either up or down?

OpenStudy (anonymous):

yes, the shape of the curve remains unchanged, it just goes up or down and don't call it a "line" or your math teacher will think you are confused

OpenStudy (anonymous):

Thanks, i just put that is shift the graph of that equation up or down

OpenStudy (anonymous):

k sound good done?

OpenStudy (anonymous):

Done, thanks so much!

OpenStudy (anonymous):

yw you owe me a beer

OpenStudy (anonymous):

haha i don't know how you didn't give up.

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

oh i never give up you did very well, especially with the one about the boat i think you got most of this, even if you are a bit confused about the log

OpenStudy (anonymous):

I have another study guide for this class for extra credit but its a lot so idk if im going to do it tonight or not.

OpenStudy (anonymous):

take a break

OpenStudy (anonymous):

I am, i might complete it in a bit, but as of now im going to relax.

OpenStudy (anonymous):

me, i am going to go to bed later

OpenStudy (anonymous):

Bye :) Thanks again.